Please how do I solve this problem: (1) 3x + 7y = 9
(2) y = 11/7 - 3x/7

Answer by ikleyn(52787)

.
Please how do I solve this problem: (1) 3x + 7y = 9
(2) y = 11/7 - 3x/7
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

3x + 7y = 9,      (1)
y = 11/7 - 3x/7.  (2)


Multiply equation (2) by 7 (both sides). You will get 

7y = 11 - 3x,     (3)

or

3x + 7y = 11.     (4)


Now you have the system of two equations

3x + 7y =  9      (1)
3x + 7y = 11      (4)

Left sides are identical, but the right sides are different.

What does it mean?   - Correct, the system has no solutions.